An electron collides elastically with a stationary hydrogen atom. The mass of the hydrogen atom is 1837 times that of the electron. Assume that all motion, before and after the collision, occurs along the same straight line. What is the final velocity of the hydrogen atom after the collision if the final velocity of the electron is -155 m/s?
Answer #1

we need to find initial velocity by using expression

v1 = (m1 - m2) u1 / ( m2 + m2)

- 155 = ( m - 1837 m) u1 / ( m +1837m)

u1 = 155.169 m/s

velocity of Hydrogen atom

v = 2 m u1 / ( m + 1837m)

v = 2* 155.169 / 1838

v = 0.16885 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck